OpenStudy (anonymous):

A car travels along a stretch of road. It proceeds for 15.9 mi at 51 mi/h, then 21.4 mi at 46 mi/h, and finally 36.8 mi at 38.6 mi/h. What's the car's average velocity during the entire trip?

OpenStudy (anonymous):

Since average velocity is total time divided by total distance, first add all of the distances together (ie 15.9 mi, 21.4 mi, and 36.8 mi) to find the total distance traveled. Then you need to find the time he took to travel each of those distances, so you need to calculate the time for each of the three portions of the trip separately. To calculate the time for any part of the trip, divide the distance by the speed (from the equation v = distance/time, solving for time). Then add all of these times together and divide by the distance you calculated earlier. Hope this helps!

OpenStudy (anonymous):

average velocity of the car = total distance covered / total time total distance = 15.9 + 21.4+ 36.8= 74.1 time 1 = 15.9/51 time 2 = 21.4/46 time 3 = 36.8/38.6 total time = 0.3117 + 0.4652 + 0.9533 so average speed = 74.1 / 1.7302 =42.827 mi / h

OpenStudy (anonymous):

thank y'all!:)